OpenStudy (anonymous):

A rectangle has a perimeter of 34 cm2 and an area of 52cm2 its length is 5 more than twice its width write and solve a system of equations to find the dimensions of the rectangle

OpenStudy (anonymous):

please help will give medal.

OpenStudy (anonymous):

@eric_d

OpenStudy (hockeychick23):

p= 34 A= 52 cm L=2w+5 w+w+2w+5+2w+5=P So now try and add up all the w's and all the numbers and see what you get

OpenStudy (anonymous):

thank you. it is 6w+10=p?

OpenStudy (hockeychick23):

correct :) and you know that p is 34 so you can plug that in: 34= 6w+10 now subtract 10 from both sides

OpenStudy (anonymous):

now were just doing simple algebraic equations 24=6w and that makes it be 4=w

OpenStudy (hockeychick23):

correct!(:

OpenStudy (anonymous):

so how would it be answered? the answer would be the unsimplified equation of 6w+10=p? I just had trouble with the wording it really conufsed me

OpenStudy (anonymous):

or would the answer be 4=w?

OpenStudy (hockeychick23):

ok so since its asking for all the dimensions and you know that 1 width=4 and you have 2 widths that would total 8, so you subtract 34-8= 26, and you have to split that between 2 lengths so divided by 2, each length is 13 and each width is 4, those are your dimensions

OpenStudy (anonymous):

but the answer is just the equation it doesnt need to be answered, at leas the way i read it, from what i read I just need to know which equation it the answer

OpenStudy (hockeychick23):

ok what are your choices?

OpenStudy (anonymous):

nvm it does need to be solved

OpenStudy (hockeychick23):

ok

OpenStudy (anonymous):

so um. I really am unsure what it is.

OpenStudy (anonymous):

hello?

OpenStudy (hockeychick23):

what are you trying to figure out? i thought you wanted the dimensions